FERMI Machine Status

Status

Run 32
Shift Machine Preparation
Beam to Gun
Next users dedicated injection 2018-02-01
Last users dedicated injection
Control Room Fermi (tel. 8800)
Info updated at

Linac / e-beam

RF plants ON
E-beam Energy 0 GeV
Bunch Charge 48 pC

Seed laser

Status OPEN
SLU Status laser OFF
Pulse Energy 35 µJ
Wavelength 265.86 nm

FEL-2 I0 monitor charts (30 minutes)

FEL-2

1st stage 2nd stage
Harmonic 10 3
Wavelength 26.59 nm 5.42 nm
Polarization Right Circular Right Circular

FEL-2 I0 monitor

1st stage 2nd stage
Intensity -0.1 µJ -16.1 µJ
RMS 35 % 5 %

Spectrometer

Wavelength 26.43 nm
Bandwidth 0.842 nm
Last Updated on Friday, 21 August 2015 15:46